gainard's grounds, a new coffee shop, mixes two types of coffe in their signature brew. the first type is a blend consisting of 7/10 kona beans. the second type consists of only 2/5 kona beans. how much of each type need to be mixed in order to produce a 50 ounce blend containing 1/2 kona beans
Answer by stanbon(75887)

gainard's grounds, a new coffee shop, mixes two types of coffe in their signature brew. the first type is a blend consisting of 7/10 kona beans. the second type consists of only 2/5 kona beans. how much of each type need to be mixed in order to produce a 50 ounce blend containing 1/2 kona
-------------
Equation:
kona + kona = kona
0.7x + (2/5)(50-x) = (1/2)*50
---
Multiply thru by 10 to get:
7x + 4*50-4*x = 5*50
---
3x = 50
x = 16 2/3 ounces (amt. of 70% needed)
50- 16 2/3 = 33 1/3 ounces (amt. of 40% needed)
==========================================
Cheers,
Stan H.
============